Write down formula for energy in wkb method when potential is x^n.
1
Expert's answer
2020-10-30T13:20:05-0400

Solution

If potential is n power of x or

"V=A x^n"

Then it's energy by using WKB method written as

"E=A\times\\((N-\frac{1}{2}) \hbar\sqrt\frac{\pi}{2m\alpha}\frac{\top(\frac{1}{n}+\frac{3}{2}) }{\top(\frac{1}{n}+1)})^{\frac{2n}{2+n}}"